# These $\equalcontThese authors contributed equally to this work.$

### v2

@@ 3.2 Sparse Gradients in Training with CE loss
Figure 4: Histograms of the gradient distributions at the final layer at the beginning (left) and end (right) of the training, the Megamarket dataset.
3.3 Cut-Cross-Entropy with Negative Sampling
As demonstrated in [16], negative sampling offers dual advantages: it allows better accuracy on certain datasets and accelerates training compared to full vocabulary utilization. Motivated by this, we extend $\mathrm{CCE}^{-}$ by integrating negative sampling into it.
$\mathrm{CCE}^{-}$ Loss Formulation. The loss function operates on an embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, which is obtained by flattening the 3D output tensor along its first two dimensions. During training, at each time step, the positive items corresponding to users are augmented with a dynamically sampled set of $ns$ negative items. This creates a compact item subset $V_{s}\subset V$ where: $|V_{s}|=1+ns$ (1 positive + $ns$ negatives) and $|V_{s}|\ll|V|$ (sampled subset size $\ll$ size of the full catalog of items ). The loss calculation is thereby confined to interactions within this compact subset, enabling efficient training, as discussed below. The $\mathrm{CCE}^{-}$loss is written as
$\mathcal{L}_{NS}=-\dfrac{1}{N}\sum_{i=1}^{N}\mathbf{l}^{NS}_{i},$ (3)
$\mathbf{l}^{NS}=(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}-\mathrm{log}\sum_{j=1}^{V_{s}}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}_{j}}}.$ (4)
@@ 3.3 Cut-Cross-Entropy with Negative Sampling
$LSE=\mathrm{log}\sum_{j=1}^{V_{s}}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}_{j}}},$ (5)
@@
To achieve a memory-efficient computation for this term, we developed a specialized kernel that fuses these operations. For this purpose, we adopt an approach for the online softmax computation. As a result, all intermediate operations are carried out on-chip using shared memory, while only the log-sum-exp vector is stored in the global memory [23, 12].
@@
$\nabla E_{i}=\sum_{j=1}^{|V_{s}|}(\mathbf{S}\cdot\nabla LSE)_{ij}C_{x_{ij}},\quad\nabla C_{x_{ij}}^{\mathrm{T}}=(\mathbf{S}\cdot\nabla LSE)_{j}^{\mathrm{T}}\cdot E_{i},$ (6)
@@
During the forward pass, the output embedding $\mathbf{E}$ is divided into the blocks $E_{n,D}$ of size $N_{B}\times D$. Since the inner dimension $D$ in transformer models for RecSys is typically not large (in  [15], $D<50$; in  [16][17], $D=64$; in  [9], $D<400$ ), division is not performed along the $D$ dimension. The matrix of the linear classifier $\mathbf{C}$ is loaded inside a for-loop by columns, determined by the block of indices $\mathbf{Inds}_{n,j}$ of size $N_{B}$.
@@
4 Experimental Settings
4.1 Datasets
Table 2: Statistics for used datasets after preprocessing
@@ 4.1 Datasets
We use six datasets: 30Music  [33], Zvuk  [26], MovieLens-20M (ML 20M)  [11], MegaMarket [26], Beauty  [20], and Gowalla  [6]. Our main focus is on datasets with a large item catalog: Zvuk, MegaMarket, Gowalla, and 30Music, which include more than 300K items. Moreover, Zvuk and MegaMarket are some of the largest open datasets in the recommendation domain  [26]. MovieLens and Beauty are added because of their popularity, used in more than 14 and 45 papers according to [17]. For Beauty, we employ its third version, which has the largest item catalog. The characteristics of all datasets after preprocessing are in Table 2.
4.2 Evaluation Methodology
@@
Global Temporal Cutoff: To ensure alignment with real-world sequential prediction scenarios, we use a global temporal split with the last interaction as a target and ”by user” validation scheme [10]. We split interactions at the 90% timestamp quantile. All interactions before this cutoff form the training/validation pool, while interactions after constitute the test set. Validation Set: From the training pool, we randomly select 5% of users. For each selected user, their last interaction is withheld for validation. Test Set: For users in the post-cutoff test set, only their last interaction is retained for evaluation, while all previous ones are passed as model input.
@@
Following prior work  [2, 40, 26], we prioritize NDCG@10 due to its strong correlation with overall ranking quality. Along with it, record other metrics, including Coverage and Surprisal [34]. All metrics are available in our publicly released dataset.
@@
We primarily focus on the SASRec model, as it is more commonly used in practice and frequently outperforms Bert4Rec [16, 21]. However, to demonstrate the efficiency of CCE and $\mathrm{CCE}^{-}$ for other architectures, we also present the results of applying these loss functions to the training of BERT4Rec in Section 5.6
@@ 4.3 Training Details
For all datasets, we trained models on a single Nvidia A100 GPU. Prior to our experiments, we tuned the number of self-attention blocks, attention heads, and hidden size. As a result, the model configuration was set to two self-attention blocks with two attention heads and a hidden size $D$ of $256$. The models were trained using the Adam optimizer with a learning rate of 1e31e-3. The training was performed with mixed-precision (float16) using the PyTorch Lightning trainer [22].
5 Results
@@
In the following, we present the results of our computational experiments and analyze them to address the questions Q1 and Q2 posed in Section 1 and evaluate the performance of CCE and $\mathrm{CCE}^{-}$. Specifically, (1) Section 5.1 contains the results of experiments carried out to study the effect of training hyperparameters on SASRec performance, where, among other insights, we demonstrate that sometimes increasing $ns$ does not lead to an improvement in the accuracy of the model; therefore, sometimes $\mathrm{CE}^{-}$ should be preferred over CE. (2) Further, we compare the training efficiency of the SASRec model using CCE and $\mathrm{CCE}^{-}$ over CE, $\mathrm{CE}^{-}$ in Sections 5.2. The results of this section further highlight the effectiveness of negative sampling in SRS. (3) Since uniform negative sampling is not the only possible sampling strategy in Section 5.4.1, we study other negative sampling strategies. (4) Finally, we extend our empirical analysis to the Bert4Rec model in Section 5.6.
5.1 Optimal Accuracy-Memory Trade-off
@@
Figure 5: Memory scaling across three dimensions: (1) marker size (slsl: sequence length), (2) color (bsbs: batch size), and (3) marker shape (nsns: number of negative samples). Asterisks denote points using the CE loss. For example, brown markers (larger batch sizes) generally yield better average performance, though the largest batch size does not consistently dominate. Higher NDCG occurs with 128–512 negative samples (squares-hexagons); only the Zvuk dataset shows improved performance with asterisks (full $\mathrm{CE}^{-}$ loss).
@@
Figure 6: This figure presents aggregated results from Figure 5, averaging across sequence lengths. Error bars indicate standard deviations. Colors represent batch sizes ($bs$). Performance increases with negative sample size ($ns$) for some datasets, while saturating for others (Movielens-20M, Gowalla). Batch size scaling shows no universal pattern: larger batches generally improve performance but saturate for 30Music, while smaller batches (128 vs 512) yield better results for Movielens-20M
@@
The first research question is how to determine the optimal configuration of training hyperparameters given the memory constraints and how specific memory-related hyperparameters affect performance. We investigate how batch size (bs), the number of negatives (ns), and sequence length (sl) affect the NDCG@10 metric of the SASRec model after training. We used a fixed grid of configurations, evaluating approximately $170$ points per dataset, resulting in a total of $997$ runs across all datasets. While the grid does not provide a fully uniform coverage of the hyperparameter space, the analysis can still be interpreted under the assumption of reasonably even exploration [25, 28].
@@
Figure 7: Parameter analysis: (a) correlation structure, (b) predictive importance for recommendation metrics, nsbs,nssl,slbsns-bs,ns-sl,sl-bs denote pairwise products as dependent variables.
Table 3: Comparison of CE and CCE. Hyperparameters such as $bs,sl$, and $ns$ were chosen based on our initial pool of scaling experiments for CE and $\mathrm{CE}^{-}$, respectively. CCE and $\mathrm{CCE}^{-}$ were trained using the same settings. We report NDCG@10, allocated memory (GB) and time per epoch (seconds). Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to CE and $\mathrm{CE}^{-}$, green denotes CCE and $\mathrm{CCE}^{-}$ superiority.
@@
(bs/sl) NDCG@10 Mem, GB Time, s NDCG@10 Mem, GB Time, s
@@
(bs/sl/ns) NDCG@10 Mem, GB Time, s NDCG@10 Mem, GB Time, s
@@
Table 4: Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except $ns$ for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(bs/sl/ns) NDCG Mem, Time, NDCG Mem, Time, % % %
@@
Table 5: Comparison Coverage@10 and Surprisal@10 time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except $ns$ for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(bs/sl/ns) Coverage Surprisal Coverage Surprisal % %
@@
Table 6: Results for $\mathrm{CCE}^{-}$ with popularity sampling on the SasRec model using identical hyperparameters as for global uniform sampling
@@
Table 7: Results for the comparison of $\mathrm{CCE}^{-}$ loss with global uniform and popularity sampling for the SASRec model. $\Delta$: Relative difference compared to popularity sampling and global uniform sampling, green denotes global uniform sampling superiority.
@@
5.3 Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
The results discussed in the previous section indicate that CCE and $\mathrm{CCE}^{-}$ significantly reduce the memory consumption required for training SASRec. This reduction provides an opportunity to enhance model performance by increasing batch size ($bs$), sequence length ($sl$), and number of samples ($ns$) in the case of $\mathrm{CCE}^{-}$, implementing insights from the Section 5.1. In these experiments, we set identical training hyperparameters for both CCE and $\mathrm{CCE}^{-}$. Tables 4, 5 present the values of the NDSG@10, Coverage@10, and Surprisal@10 metrics for all datasets, along with the allocated memory and time per epoch required for training.
@@ 5.3 Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
5.4 Gradient Filtering with CCE
@@
Figure 8: Performance of SASRec trained with CCE using gradient filtering at increasing thresholds (filter_eps parameter). The NDCG@10 metric (left axis) and the training time per epoch (right axis) are shown.
@@
Figure 9: NDCG@10 metric and time per epoch of the SASRec models trained with $\mathrm{CCE}^{-}$ and CCE using gradient filtering at increasing thresholds.
5.4.1 Comparison Sampling Strategies

### v3

@@ 3.2. Sparse Gradients in Training with CE loss
Figure 4. Histograms of the gradient distributions at the final layer at the beginning (left) and end (right) of the training, the Megamarket dataset.
$3.$ Cut-Cross-Entropy with Negative Sampling
As demonstrated in (klenitskiy2023turning), negative sampling offers dual advantages: it allows better accuracy on certain datasets and accelerates training compared to full vocabulary utilization. Motivated by this, we extend $\mathrm{CCE}^{-}$ by integrating negative sampling into it.
$\mathrm{CCE}^{-}$ Loss Formulation. The loss function operates on an embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$, which is obtained by flattening the 3D output tensor along its first two dimensions. During training, at each time step, the positive items corresponding to users are augmented with a dynamically sampled set of nsns negative items. This creates a compact item subset $V_{s}\subset V$ where: $|V_{s}|=1+ns$ (1 positive + nsns negatives) and $|V_{s}|\ll|V|$ (sampled subset size $\ll$ size of the full catalog of items ). The loss calculation is thereby confined to interactions within this compact subset, enabling efficient training, as discussed below. The $\mathrm{CCE}^{-}$loss is written as
(3) $\mathcal{L}_{NS}=-\dfrac{1}{N}\sum_{i=1}^{N}\mathbf{l}^{NS}_{i},$
(4) $\mathbf{l}^{NS}=(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{x}}-\mathrm{log}\sum_{j=1}^{V_{s}}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}_{j}}}.$
@@
(5) $LSE=\mathrm{log}\sum_{j=1}^{V_{s}}\exp{(\mathbf{C}^{\mathrm{T}}\mathbf{E})_{\mathbf{Inds}_{j}}},$
@@
To achieve a memory-efficient computation for this term, we developed a specialized kernel that fuses these operations. For this purpose, we adopt an approach for the online softmax computation. As a result, all intermediate operations are carried out on-chip using shared memory, while only the log-sum-exp vector is stored in the global memory (milakov2018onlinesoftmax; hsuliger).
@@
(6) $\nabla E_{i}=\sum_{j=1}^{|V_{s}|}(\mathbf{S}\cdot\nabla LSE)_{ij}C_{x_{ij}},\quad\nabla C_{x_{ij}}^{\mathrm{T}}=(\mathbf{S}\cdot\nabla LSE)_{j}^{\mathrm{T}}\cdot E_{i},$
@@
During the forward pass, the output embedding $\mathbf{E}$ is divided into the blocks $E_{n,D}$ of size $N_{B}\times D$. Since the inner dimension $D$ in transformer models for RecSys is typically not large (in  (kang2018self), $D<50$; in  (klenitskiy2023turning)(klenitskiy2024does), $D=64$; in  (guo2024scaling), $D<400$ ), division is not performed along the $D$ dimension. The matrix of the linear classifier $\mathbf{C}$ is loaded inside a for-loop by columns, determined by the block of indices $\mathbf{Inds}_{n,j}$ of size $N_{B}$.
@@
4. Experimental Settings
4.1. Datasets
Table 2. Statistics for used datasets after preprocessing
@@ 4.1. Datasets
We use six datasets: 30Music  (turrin201530music), Zvuk  (shevchenko2024variability), MovieLens-20M (ML 20M)  (harper2015movielens), MegaMarket (shevchenko2024variability), Beauty  (mcauley2015image), and Gowalla  (cheng2013you). Our main focus is on datasets with a large item catalog: Zvuk, MegaMarket, Gowalla, and 30Music, which include more than 300K items. Moreover, Zvuk and MegaMarket are some of the largest open datasets in the recommendation domain  (shevchenko2024variability). MovieLens and Beauty are added because of their popularity, used in more than 14 and 45 papers according to (klenitskiy2024does). For Beauty, we employ its third version, which has the largest item catalog. The characteristics of all datasets after preprocessing are in Table 2.
4.2. Evaluation Methodology
@@
Global Temporal Cutoff: To ensure alignment with real-world sequential prediction scenarios, we use a global temporal split with the last interaction as a target and ”by user” validation scheme (gusak2025time). We split interactions at the 90% timestamp quantile. All interactions before this cutoff form the training/validation pool, while interactions after constitute the test set. Validation Set: From the training pool, we randomly select 5% of users. For each selected user, their last interaction is withheld for validation. Test Set: For users in the post-cutoff test set, only their last interaction is retained for evaluation, while all previous ones are passed as model input.
@@
Following prior work  (anelli2019discriminative; zhao2022revisiting; shevchenko2024variability), we prioritize NDCG@10 due to its strong correlation with overall ranking quality. Along with it, record other metrics, including Coverage and Surprisal (vasilev2024replay). All metrics are available in our publicly released dataset.
@@
We primarily focus on the SASRec model, as it is more commonly used in practice and frequently outperforms Bert4Rec (klenitskiy2023turning; mezentsev2024scalable). However, to demonstrate the efficiency of CCE and $\mathrm{CCE}^{-}$ for other architectures, we also present the results of applying these loss functions to the training of BERT4Rec in Section 5.6
@@ 4.3. Training Details
For all datasets, we trained models on a single Nvidia A100 GPU. Prior to our experiments, we tuned the number of self-attention blocks, attention heads, and hidden size. As a result, the model configuration was set to two self-attention blocks with two attention heads and a hidden size $D$ of $256$. The models were trained using the Adam optimizer with a learning rate of 1e31e-3. The training was performed with mixed-precision (float16) using the PyTorch Lightning trainer (micikevicius2017mixed).
5. Results
@@
In the following, we present the results of our computational experiments and analyze them to address the questions Q1 and Q2 posed in Section 1 and evaluate the performance of CCE and $\mathrm{CCE}^{-}$. Specifically, (1) Section 5.1 contains the results of experiments carried out to study the effect of training hyperparameters on SASRec performance, where, among other insights, we demonstrate that sometimes increasing nsns does not lead to an improvement in the accuracy of the model; therefore, sometimes $\mathrm{CE}^{-}$ should be preferred over CE. (2) Further, we compare the training efficiency of the SASRec model using CCE and $\mathrm{CCE}^{-}$ over CE, $\mathrm{CE}^{-}$ in Sections 5.2. The results of this section further highlight the effectiveness of negative sampling in SRS. (3) Since uniform negative sampling is not the only possible sampling strategy in Section 5.4.1, we study other negative sampling strategies. (4) Finally, we extend our empirical analysis to the Bert4Rec model in Section 5.6.
5.1. Optimal Accuracy-Memory Trade-off
@@
Figure 5. Memory scaling across three dimensions: (1) marker size (slsl: sequence length), (2) color (bsbs: batch size), and (3) marker shape (nsns: number of negative samples). Asterisks denote points using the CE loss. For example, brown markers (larger batch sizes) generally yield better average performance, though the largest batch size does not consistently dominate. Higher NDCG occurs with 128–512 negative samples (squares-hexagons); only the Zvuk dataset shows improved performance with asterisks (full $\mathrm{CE}^{-}$ loss).
@@
Figure 6. This figure presents aggregated results from Figure 5, averaging across sequence lengths. Error bars indicate standard deviations. Colors represent batch sizes (bsbs). Performance increases with negative sample size (nsns) for some datasets, while saturating for others (Movielens-20M, Gowalla). Batch size scaling shows no universal pattern: larger batches generally improve performance but saturate for 30Music, while smaller batches (128 vs 512) yield better results for Movielens-20M
@@
The first research question is how to determine the optimal configuration of training hyperparameters given the memory constraints and how specific memory-related hyperparameters affect performance. We investigate how batch size (bs), the number of negatives (ns), and sequence length (sl) affect the NDCG@10 metric of the SASRec model after training. We used a fixed grid of configurations, evaluating approximately $170$ points per dataset, resulting in a total of $997$ runs across all datasets. While the grid does not provide a fully uniform coverage of the hyperparameter space, the analysis can still be interpreted under the assumption of reasonably even exploration (scarlett2017lower; srinivas2010gaussian).
@@
Figure 7. Parameter analysis: (a) correlation structure, (b) predictive importance for recommendation metrics, nsbs,nssl,slbsns-bs,ns-sl,sl-bs denote pairwise products as dependent variables.
Table 3. Comparison of CE and CCE. Hyperparameters such as $bs,sl$, and nsns were chosen based on our initial pool of scaling experiments for CE and $\mathrm{CE}^{-}$, respectively. CCE and $\mathrm{CCE}^{-}$ were trained using the same settings. We report NDCG@10, allocated memory (GB) and time per epoch (seconds). Bold values indicate best performance per dataset. $\Delta$: Relative difference compared to CE and $\mathrm{CE}^{-}$, green denotes CCE and $\mathrm{CCE}^{-}$ superiority.
@@
(lr)3-5 (lr)6-8 (bs/sl) NDCG@10 Mem, GB Time, s NDCG@10 Mem, GB Time, s
@@
(lr)3-5 (lr)6-8 (bs/sl/ns) NDCG@10 Mem, GB Time, s NDCG@10 Mem, GB Time, s
@@
Table 4. Comparison NDCG@10, memory consumption and epoch time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except nsns for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(lr)3-5 (lr)6-8 (bs/sl/ns) NDCG Mem, Time, NDCG Mem, Time, % % %
@@
Table 5. Comparison Coverage@10 and Surprisal@10 time for CCE and $\mathrm{CCE}^{-}$ on the SASRec model using identical training hyperparameters, except nsns for $\mathrm{CCE}^{-}$. Bold values indicate the best performance per dataset. $\Delta$: Relative difference compared to $\mathrm{CCE}^{-}$, green denotes $\mathrm{CCE}^{-}$ superiority.
@@
(lr)3-4 (lr)5-6 (lr)7-7 (lr)8-8 (bs/sl/ns) Coverage Surprisal Coverage Surprisal % %
@@
Table 6. Results for $\mathrm{CCE}^{-}$ with popularity sampling on the SasRec model using identical hyperparameters as for global uniform sampling
@@
Table 7. Results for the comparison of $\mathrm{CCE}^{-}$ loss with global uniform and popularity sampling for the SASRec model. $\Delta$: Relative difference compared to popularity sampling and global uniform sampling, green denotes global uniform sampling superiority.
@@
5.3. Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
The results discussed in the previous section indicate that CCE and $\mathrm{CCE}^{-}$ significantly reduce the memory consumption required for training SASRec. This reduction provides an opportunity to enhance model performance by increasing batch size (bsbs), sequence length (slsl), and number of samples (nsns) in the case of $\mathrm{CCE}^{-}$, implementing insights from the Section 5.1. In these experiments, we set identical training hyperparameters for both CCE and $\mathrm{CCE}^{-}$. Tables 4, 5 present the values of the NDSG@10, Coverage@10, and Surprisal@10 metrics for all datasets, along with the allocated memory and time per epoch required for training.
@@ 5.3. Reallocating memory savings to increase performance with CCE and $\mathrm{CCE}^{-}$
5.4. Gradient Filtering with CCE
@@
Figure 8. Performance of SASRec trained with CCE using gradient filtering at increasing thresholds (filter_eps parameter). The NDCG@10 metric (left axis) and the training time per epoch (right axis) are shown.
@@
Figure 9. NDCG@10 metric and time per epoch of the SASRec models trained with $\mathrm{CCE}^{-}$ and CCE using gradient filtering at increasing thresholds.
5.4.1. Comparison Sampling Strategies